OpenStudy (anonymous):

What is the missing number in this sequence? 15 , 18 , 24 , ___ , 45

OpenStudy (anonymous):

A 36 B 31 C 33

OpenStudy (anonymous):

What do you think it is? :P

OpenStudy (anonymous):

B

OpenStudy (anonymous):

That's the least likely one, actually do it instead of guessing, because us giving you answers isn't helping you :P

OpenStudy (whpalmer4):

15 + 3 = 18 18 + 6 = 24 24 + ??? = ????? ????? + ????? = 45

OpenStudy (whpalmer4):

What do you think a reasonable possibility for the number ??? in 24 + ??? might be, if you follow the pattern established by the first two lines?

OpenStudy (anonymous):

Katie do you see a pattern in what @whpalmer4 wrote?

OpenStudy (whpalmer4):

Or, looking at it from a different angle, if you think that 31 is the number that goes in the sequence: 15, 18, 24, 31, 45 What is the rule that you use to find the next number?

OpenStudy (anonymous):

is it 36

OpenStudy (whpalmer4):

I'm not going to play the guessing game with you. You'll need to figure out the rule. You should be able to do so.

OpenStudy (anonymous):

I agree, I've gave you too many answers as it is, which I shouldn't have :X

OpenStudy (whpalmer4):

If you think the sequence is 15, 18, 24, 36, 45 What is the rule that constructs the next number?

OpenStudy (anonymous):

im not guessing I have a learning problem

OpenStudy (anonymous):

what's 15 + 3 =?

OpenStudy (anonymous):

18...

OpenStudy (anonymous):

ok what's 18 + 6 =?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

24 + ? = ? Do you see the pattern yet with the numbers you have to add?

OpenStudy (anonymous):

From 15 to 18, you added three. From 18 to 24, you added six. To get from 24 to the next number, how much do you need to add? And from that number to 45, you will need to add...? Does this work?

OpenStudy (whpalmer4):

So far we've added 3 to the first number to get the second number we've added 6 to the second number to get the third number what might we reasonably add to the third number to get the fourth number? Seems like two obvious choices would be 9 (adding 3 more each time), or 12 (doubling each time). How to tell?

OpenStudy (anonymous):